Sara likes consume Bananas (X) and Mangoes (Y). The cost of banana is birr 2 and that of Mango is bírr 4. Sara's monthly budget for these fruit is birr 100 and her utility function is U(X,Y) = X ^ 2 * Y ^ 1/2,Then calculate Utility maximizing units of Bananas and Mangoes.

1
Expert's answer
2022-03-24T16:12:35-0400

U(xy)=X"^2" Y0.5

Budget line,where px=2,py=4

PxX+PyY=100

2X+4Y=100

Utility is maximized when

"\frac{MUx}{MUy}=\frac{Px}{Py}"

MUx="\frac{du}{dy}" X"^2" Y0.5


MUx=2XY0.5


MUy=0.5X"^2"Y-0.5

therefore,

"\frac{MUx}{MUy}=\frac{Px}{Py}"


"\frac{2XY^\frac{1}{2}}{0.5X^2Y^\frac{-1}{2}}=\frac{2}{4}"


"\frac{2y}{0.5x}=\frac{1}{2}" ,


0.5x=4y

x=8y

substitute in budget line function

2x+4y=100

2(8y)+4y=100

16y+4y=100

20y=100

y=5


2x+4y=100

2x+(4"\times5)" =100

2x+20=100

2x=100-20

2x=80

x=40

maximum utility Bananas(x)=40,Mangoes(y)=5